OpenStudy (anonymous):

HELP! :) What is the 9th term of the geometric sequence where a1 = 1024 and a3 = 64?

OpenStudy (anonymous):

0.015625 0.0625 0.125 0.25

OpenStudy (anonymous):

I think its A? :)

OpenStudy (anonymous):

geometric seequence means : \[a_n=q^{n-1}a_1\] so \[\frac{a_3}{a_1}=q^2=\frac{64}{1024}=1/16\] so q=1/4 now: \[a_9=(1/4)^81024=0.015625\]